Here is a question we could ask person n. For what magnitude of are you indifferent between having 5k for sure and having 1k with probability T1 and 10k with probability (1-m). Suppose person n answers π1-3. If we interpret that answer in terms of expected utility, then we have the equation where the second equality comes from the normalization, un (1k)-0 and un (10k)-1 Having determined that un(5k) - .7, we could ask other questions. Foir example, we could ask, for what magnitude of π1 is person n indifferent between having 3k for sure and having 1k with probability Ti and 5k with probability (1-T). Suppose they answer. π1-2. If we interpret that answer in terms of expected utility, then we have the equation where the second equality comes from the normalization, un(1k) = 0 and the previous result that un(5k) = .7. Notice that we are eliciting person ns attitude toward risk. There are no ri ght or wrong answers to these answers Exercise 2 Put yourself in the position of person n and pose and answer 4 questions that permit you to find your u function for 4 different magnitudes of r between 1k and 10k. Use some software to sketch those points of your u function.
Answer #1

Let x be the monetary value in $1000. Let u(x) indicate the utility of having $x (in thousands) (Here we are assuming that the monetary values are dollars although the question does not specify it)

We will first set the 2 values of the utilities at x=1 as u(x=1)=0 and u(x=10)=1. That is utility of $1k is set at 0 and utility of $10k is set at 1.

We want to fill in the following table, by eliciting responses for x=2,4,6, and 8 (in thousand dollars)

x ($1000) u(x)
1 0
2
4
6
8
10 1
  • Question 1: What value of pi_1 that you are indifferent to receiving $4k for sure vs a lottery that will pay $1k with probability pi_1 and $10k with probability (1-pi_1). A risk seeker will go for a higher value of pi_1 (riskier bet) and a risk averse person will go for a lower value of pi_1 (lower risk) (higher value is towards 1 and lower value is toward 0)
    • ans: pi_1=0.4 (You can change these values, but you get what we are trying to achieve)

u(z = 4) = π1u(1) + (1-m)u(10) = 0.4 . O+ (1-0.4) × 1 = 0.6

Hence the utility of $4k for me is 0.60

  • Question 2: What value of pi_1 that you are indifferent to receiving $2k for sure vs a lottery that will pay $1k with probability pi_1 and $4k with probability (1-pi_1).
    • ans: = 0.5

u(z. 2)--111 (1) + (1--1 )11(4) = 0.5 × 0 + (1-0.5) × 0.6-0.3

Hence the utility of $2k for me is 0.30

  • Question 3: What value of pi_1 that you are indifferent to receiving $6k for sure vs a lottery that will pay $4k with probability pi_1 and $10k with probability (1-pi_1).
    • ans: pi_1=0.6

u(z = 6) = π1u(4) + (1-m)a(10) = 0.6 × 0.6 + (1-0.6) × 1 = 0.76

Hence the utility of $6k for me is 0.76

  • Question 4: What value of pi_1 that you are indifferent to receiving $8k for sure vs a lottery that will pay $6k with probability pi_1 and $10k with probability (1-pi_1).
    • ans: pi_1=0.65

u(z 8) = π1a(6) + (1-m)a(10) = 0.65 × 076 + (1-0.65) × 1 = 0.844

Hence the utility of $8k for me is 0.844

We get the following table of utilities

x ($1000) u(x)
1 0
2 0.3
4 0.6
6 0.76
8 0.844
10 1

We plot this in excel using insert-->scatter--->lines

File Home InsrtPage Layout Formulas Data Review VieDeveloper Add-Ins PowerPivot PivotTable Table Picture Clip Shapes SmartArt Screenshot Column Line Pe AaScatter OtherLi Column Win/Loss Art Chart Tables Illustrations Charts Scatter parklines bella! allles KAI I Compare pairs of values. A1 E x ($1000) 1 x ($1000) u(x) 0 0.3 0.6 6 0.76 8 0.844 2 4 d. Scatter with Straight Lines and M Use it when there are a few dat All Chard Points in x-axis order and the d represents separate values

get the following raw graph

plot of u(x) 1.2 ︵ 0.8 0.6 0.4 0.2 0 4 10 x (thousands)

Since this utility is largely concave, this is a utility function for a risk averse person. But it turns up at higher payoffs, indicating a slightly higher risk apatite at higher payoffs.